# AI for Just Work: Constructing Diverse Imaginations of AI beyond “Replacing Humans”

### v1

@@ 3 Ethical and diverse imaginations are needed for AI
In the previous section, we critically examine the MIA and defamiliarize ordinary consensus that we take for granted in the community. People may defend MIA by the benefits brought by AI. Our examination does not reject the benefits either. The disagreement lies, is it really necessary to couple AI development with the harms? Or, is it possible to imagine different visions of developing AI that have the benefits and get rid of the harms? Our critical examination points out that the root causes of harms are not in technology itself, but lie in the flawed assumptions in the imagination that underpins technology. These flawed assumptions are not a necessary component in technological development; rather, they are highly dependent on our choices of how to shape our imagination. Technological development is not deterministic, and harms may not be an inevitable side effect of technology. This means that it is possible to decouple harms from AI development by getting rid of the unethical assumptions in the imaginations of AI, which is identical to replacing the unspoken and unjustified unethical assumptions with transparent and justified ethical assumptions from the outset in the imagination of AI.
@@
However, it should be noted that the reason for MIA to be the dominant imagination is that it is deeply rooted in the dominant value systems and social and power structures that set the paths of least resistance (Johnson, 2014) for people to act and succeed without the efforts to critically examine “why” questions behind actions, such as setting the agenda of trendy research topics (Rolnick et al., 2024; Rudin & Wagstaff, 2013; Wagstaff, 2012), common design assumptions, popular benchmarks (Birhane & Prabhu, 2021; Koch et al., 2021; Raji et al., 2021), and shaping common practices in evaluation (Reinke et al., 2024; Jin et al., 2024), reporting (Burnell et al., 2023), and peer review (Andrews et al., 2024; Lipton & Steinhardt, 2019). As a consequence, this greatly limited our ability to imagine organizing our social relationships and our relationship with AI in alternative visions (Benjamin, 2024; Costanza-Chock, 2023). Choosing to deviate from MIA (the path of least resistance) means that individuals and the community need to constantly investigate efforts in collective imaginations to “get over some of this deeply habituated laziness and start engaging in interpretive (imaginative) labor for a very long time to make those realities stick”  (Graeber, 2015). Despite the hard work, such endeavors are essential for the healthy development of AI and the research community, if we want to correct the current trend of AI research being captivated by private sectors’ values and interests (Young et al., 2022; Birhane et al., 2022; Zheng et al., 2024), and restore the value of AI research for public good. That means the AI community needs to be aware of the problems in MIA by engaging in scientific criticisms from inside and outside the community and making the self-correcting mechanism in scientific research work for our community, and starts to recognize, support, and cultivate diversified imaginations of AI for various ethical pursuits and enable them to be on the paths of least resistance in the research processes, such as research topic identification, funding, and peer review.
@@ 4 AI for just work: a demonstration of constructing a new imagination
Diverse imaginations can be constructed by focusing on various ethical values such as social justice, environmental justice, the values of community and solidarity, etc. And specific technical works can build upon one or the combination of several ethical imaginations. To inspire and facilitate the community to construct diversified imaginations of AI, in this section, we demonstrate our process to construct a new imagination of “AI for just work” and apply it in our specific AI development. We first construct and justify three foundational assumptions that form the value system for the new imagination, and deduce new properties of AI from these assumptions. We then show the process of how the new imagination guides specific technical decisions in the case study of medical image synthesis task.
@@ 4.1 Foundational assumptions
We define complex phenomenon by borrowing the definition of complex systems, which are “co-evolving multilayer networks,” are context-dependent, and are composed of many non-linearly interacting elements (Thurner et al., 2018), such as language, human behavior, human mind, a living cell, financial market, social or natural phenomena. A1 is justified by our counterarguments to S1 and the normal worldview of science and technology (Fig. 2-A). This worldview acknowledges the limits of science and technology that science “isn’t a window onto a disembodied, God’s-eye perspective. It doesn’t grant us access to a perfectly knowable, timeless objective reality. …Instead, all science is always our science, profoundly and irreducibly human, an expression of how we experience and interact with the world” (Frank et al., 2024). Similarly, A1 assumes that there does not exist a disembodied, God-like intelligence of knowing that grants us access to a perfect world model that can replace the original phenomena. It acknowledges and respects the fact that AI models are built upon human experience (embodied as human labor, prior knowledge, dataset, infrastructure, etc.), natural resources and materials of our planet (Crawford, 2021). This assumption is also well-supported by common framing in the academic community interested in machine learning and statistics. However, it stands in contrast to how AI is sometimes described in the tech industry.
@@
A2 is justified by our counterarguments to S2 that the root causes of AI-based algorithmic oppression and exploitation are in the social hierarchy that wrongly assumes some people and their work are inferior, and in the imbalanced power between the oppressor and the oppressed. To correct the unjust assumptions, A2 assumes that workers and their work are not devalued or degraded by any means, including but not limited to: workers’ age, gender, sexuality, race, ethnicity, religion, nationality, physical or mental conditions, educational level, socioeconomic status, whether the work is paid or unpaid, the type, role, and salary of the work, the amount of efforts or skills required or engaged, how workers are rated and their ratings, etc. Furthermore, in this new imagination, we refer to critical theory and adopt a broader definition of work to revive its original meaning as “meaningful, fulfilling activities” as the “social means necessary to live a flourishing life” (Wright, 2021). Work under this assumption includes a wider range of human activities that may not count as “work” in the current narrowly framed economic system based on market exchange, such as reproductive labor and emotional labor (Raworth, 2017; Mies, 1986). In the context of AI, workers include a wide range of people who work for or behind AI (such as data labor), who work with AI (such as users), and whose work or life is impacted by AI (such as stakeholders). Using the broad concept of work ensures that people and their visible or invisible labor and contributions can be recognized, supported, and deserve workers’ power to protect their rights. In the remaining parts of the paper, we use the word “workers” to denote the broad groups of people that include workers, users, stakeholders, and the public. The ideas in A2 are also justified by AI ethics principles of autonomy and justice (Floridi et al., 2018).
@@
A3 is justified by our counterarguments to S3 that merely pursuing productivity and efficiency cannot automatically lead to shared prosperity, as this goal can easily be hijacked by the dominant power to serve the benefits and interests of the powerful. Instead of pursuing progress and productivity at the expenses of essential ethical values, it is more reasonable to maintain dynamic balance between ethics and development for the same goal of the mutual flourishing of humanity and the planet. The ideas in A3 are also justified by AI ethics principles of beneficence and non-maleficence (Floridi et al., 2018).
@@ 4.2 AI properties
As shown in A1 and Fig. 2, the blind spot worldview (Frank et al., 2024) and the abstraction nature of AI models tend to get people caught up in the upper abstraction space, ignoring the grounded real-world space where AI models are applied and can have an impact. As Rolnick et al. (2024) point out, “ML algorithms designed in blue-sky, methods-focused research continue to fall short when used directly for applications.” According to A1, because AI models abstract and simplify real-world phenomena, the modeling process creates inevitable information loss due to the gap between concrete world and abstract model. Therefore, considerable human efforts are indispensable (visualized as pillars in Fig. 2-A) to mitigate the gap and information loss. Human effort acts in a two-way process from real world to AI model and back to real world: First, in the modeling process, human labor is needed to determine which information from the real world is relevant for AI abstraction, such as the efforts in problem formulation, data curation, data labeling, and model design; Second, to apply AI in the real world, human labor is necessary to act as a buffer between AI model and real-world tasks to deal with the limitations of model and the ambiguity and fluidity of the real world, such as efforts in model evaluation, reorganizing the existing workflow, incorporating AI outputs in real-world tasks, long-term model monitoring and quality control. The current main paradigm of methods-focused AI research is not driven by and grounded in the real world problems, making the above indispensable human/social processes not being seriously considered and incorporated in the modeling process. This greatly limits the positive impacts of AI research to society and impedes AI research itself (Rolnick et al., 2024). Furthermore, failing to ground AI models in real world applications tends to undervalue human labor in AI development and deployment, and underestimate the negative impacts of AI implementation on workers, which violates A2. P1 joins the recurring call in the AI community for real world problem-grounded AI (Rolnick et al., 2024; Rudin & Wagstaff, 2013; Wagstaff, 2012), and provides additional justifications for this call based on the new imagination.
@@
To ground the problem formulation, design, and evaluation of AI in real-world tasks, the first step is to understand and acknowledge the inevitable gap between AI modeling in the abstracted space and its application in the real world, and should not conflate the model’s capability at the abstracted level (confined by many idealized assumptions) with the model’s real-world utility and performance, which manifests the pathology of “confusing the map with the territory” in the blind spot worldview (Frank et al., 2024). The second step is trying to mitigate the gap by relaxing the too idealistic assumptions confined by the toy problem formulation and benchmarks, and incorporating more domain- and task-specific knowledge and human experience in the modeling process. As Andrews et al. (2024) state, this may improve scientific validity and avoid unethical pitfalls for AI. For example, instead of framing an evaluation problem as AI vs. doctor on a benchmark dataset, which is unrealistic in clinical scenario, the ultimate evaluation of a medical AI should assess how the AI and clinical workers perform when AI is embedded in clinical contexts (Cabitza et al., 2017).
@@
The current evaluation paradigm of AI is lopsided as it mainly evaluates the positive aspects of the proposed AI technique, such as performance improvement and efficacy of a proposed function from ablation study, but the negative aspects of the proposed AI technique, such as its limitations and negative impacts, are at most briefly mentioned without an equivalent amount of thorough analysis as the positive aspects (Herrmann et al., 2024). The biased evaluation paradigm is driven by and reinforces the MIA that technical progress of AI is to reach an idealized image of AI that is perfect with no weaknesses; and technical progress is prioritized and harms can be fixed later (S1), which violates A3 in the new imagination.
In the new imagination, we acknowledge that there does not exist a perfect state of AI, and AI always has its limitations and weaknesses no matter how technology progresses (A1). To avoid weakening workers’ power from the incomplete evaluation that only emphasizes AI strengths and human weaknesses but neglects human strengths and AI weaknesses (A2), and to avoid developing AI for the sake of technical progress and ignoring its negative aspects (A3), we propose that a thorough limitation analysis of AI should be a routine in AI evaluation. The limitation analysis includes the following aspects: 1) quantitatively and/or qualitatively assessing the scope, weaknesses, and failure modes, and associating potential social consequences with the limitations, so that users can have a holistic understanding of the strengths and risks of the proposed AI technique. 2) Being transparent and acknowledging the design-specific assumptions and limitations in problem formulation, model design, and evaluation methodology that may be mitigated in future work. 3) Acknowledging the intrinsic limits of the proposed AI technique that cannot be overcome with technological progresses. Setting up the thorough limitation analysis as an AI evaluation standard also aligns with the scientific practice in biomedicine or engineering that equally evaluate the scope, efficacy, side effects, and safety issues of a medication or an engineering product.
@@
Because AI models and applications have scopes, limitations, risks, costs, and uncertainty (A1, P2), it is reasonable to deduce that AI may not be applicable in every scenario, thus AI and its data collections can be rejected. As Benjamin (2016) states, informed refusal is the corollary of informed consent in ethics. Refusal also corresponds to the right to withdraw in research ethics, and has already been implemented in laws. For example, GDPR (2018) states “data subject shall have the right not to be subject to a decision based solely on automated processing,” and have the right to erasure data (“right to be forgotten”). These measures are meant to protect rights and shift power dynamics in favor of the powerless, in this innately imbalanced power structure where technical sectors and institutions are the dominant.
However, this refusal assumption is currently not embedded in AI technical development, which creates a strong framing effect that sets the given AI model and technical solutions as default and their necessity is rarely interrogated or challenged. As we mentioned previously in S4, not all problems are suitable to be solved by AI. Maybe the “problem” is not a real problem at all that needs to be solved, or maybe the root causes of the problem can be tackled by non-technical or low-technical approaches. Failing to recognize this framing effect falls into the “solutionism trap” (Selbst et al., 2019). Encoding refusal as the basic property of AI can avoid the trap of regarding AI and its data collection as the default and inevitable option. Furthermore, refusal is not the end goal, but a way of constructing a reciprocal relationship between the powerless and the powerful, such that the design of AI and data collection can incorporate goals and perspectives of workers, especially the vulnerable and marginalized, at the very start (Benjamin, 2016; Zong & Matias, 2024), which is in line with A2, A3, and P1.
@@
Because certain AI design choices, such as automation, can create a displacement effect that reduces workers’ wages and weakens their power (Acemoglu & Restrepo, 2019a, c), which devalues workers’ work thus violates A2, an ethical and prudent design of AI should take the impact of AI on worker rights and power into consideration, and choose to design AI that could improve, rather than weaken workers’ value in work. As we mentioned in the counterarguments to S3, such design choices can include AI that aids workers, AI that creates new labor-intensive tasks and opportunities in which workers have comparative advantages over machine (Acemoglu et al., 2023). In cases when AI displacement effect cannot be completely offset, mechanisms should be placed to compensate for the reduced workers’ rights and power, such as retraining workers for new skills. Scholars also point out that the so-called “collaborative AI,” “human-augmenting AI,” or “human-centered AI” may not necessarily lead to increased worker power, as some may still reproduce a hierarchical and exploitative relationship that shifts the oppression to the less powerful (Atanasoski & Vora, 2019a; Costanza-Chock, 2023). Therefore, whether or not an AI design reduces worker power should not be determined by the powerful party of the technical sector or the institution who implements the AI, but should undergo democratic processes by incorporating a wide range of workers’ voices and perspectives in AI design, especially the vulnerable and marginalized. Ultimately, this collective design paradigm in AI development depends on power check mechanisms and structural changes that shift power dynamics from capital to workers and tackle the root causes in social institutions that exploit workers and make work miserable, such as collective bargaining and legislation to protect worker rights and power (for instance, their rights to reject AI in P3), and changing the ownership of production from private-owned to public- or worker-owned. The AI community has the agency to propose technical works and non-technical initiatives to facilitate, rather than impede, such power checks and structural changes.

### v2

@@ 3 Ethical and diverse imaginations are needed for AI
In the previous section, we critically examine the MIA and defamiliarize ordinary consensus that we take for granted in the community. People may defend MIA by the benefits brought by AI. Our examination does not reject the benefits either. The disagreement lies, is it really necessary to couple AI development with the harms? Or, is it possible to imagine different visions of developing AI that have the benefits and get rid of the harms? Our critical examination points out that the root causes of harms are not in technology itself, but lie in the flawed assumptions in the collective imaginations that underpins technology. These flawed assumptions are not a necessary component in technological development; rather, they are highly dependent on our choices of how to shape our collective imaginations. Technological development is not deterministic, and harms may not be an inevitable side effect of technology. This means that it is possible to decouple harms from AI development by getting rid of the unethical assumptions in the imaginations of AI, which is identical to replacing the unspoken and unjustified unethical assumptions with transparent and justified ethical assumptions from the outset in the imagination of AI.
@@
However, it should be noted that the reason for MIA to be the dominant imagination is that it is deeply rooted in the dominant value systems and social and power structures that set the paths of least resistance [63] for people to act and succeed without the efforts to critically examine “why” questions behind actions, such as setting the agenda of trendy research topics [95, 97, 104], common design assumptions, popular benchmarks [21, 66, 90], and shaping common practices in evaluation [92, 61], reporting [29], and peer review [13, 71]. As a consequence, this greatly limited our ability to imagine organizing our social relationships and our relationship with AI in alternative visions [18, 37]. Choosing to deviate from MIA (the path of least resistance) means that individuals and the community need to constantly investigate efforts in collective imaginations to “get over some of this deeply habituated laziness and start engaging in interpretive (imaginative) labor for a very long time to make those realities stick”  [55]. Despite the hard work, such endeavors are essential for the healthy development of AI and the research community, if we want to correct the current trend of AI research being captivated by private sectors’ values and interests [111, 20, 112], and restore the value of AI research for public good. That means the AI community needs to be aware of the problems in MIA by engaging in scientific criticisms from inside and outside the community and making the self-correcting mechanism in scientific research work for our community, and starts to recognize, support, and cultivate diversified imaginations of AI for various ethical pursuits and enable them to be on the paths of least resistance (i.e., structural changes).
Specifically, structural changes can be made by critically scrutinizing flaws in the embedded imaginations and values in existing procedures and institutional arrangements in the AI community, such as the education, mentoring, research agenda, research assessment, peer review, funding, and career advancement systems. Improvements can be made by taking a bottom-up approach that co-designs and co-iterates new ethical and diverse imaginations embedded in the institutional arrangements with a wide range of stakeholders, especially the non-technical stakeholders who are marginalized, vulnerable, and disproportionately affected. For example, to make AI research publishing venues accountable to the public and society, venues can set up public engagement mechanisms (e.g. public meetings, consultations, advisory boards, and participatory design processes) that empower the broader public to shape the imaginations and values encoded in the research agenda and peer review criteria. Engaging with non-technical stakeholders and diverse public is vital, because flaws in the imaginations and values may not be easily recognizable by people who have homogeneous experienced and perspectives. In fact, as we mentioned, imagination intrinsically resists challenge and change, because it provides us with the deep and unconscious frame of thinking to construct a predictable and taken-for-granted reality. To make our call for diverse and ethical imaginations of AI easier to implement, next, we demonstrate our method of constructing a new imagination of AI, starting from consciously rebuilding the fundamental assumptions and then deducing concrete AI properties from them to facilitate technical implementation. Our example also shows a process of pushing for structural change toward more ethical imaginations by critically examining flaws in existing technical practice guided by the new imagination, and making ethical practice recommendations for the research community to consider and iterate with diverse stakeholders.
@@ 4 AI for just work: a demonstration of constructing a new imagination
Diverse imaginations can be constructed by focusing on various ethical values such as social justice, environmental justice, the values of community and solidarity, etc. And specific technical works can build upon one or the combination of several ethical imaginations. To inspire and facilitate the community to construct diversified imaginations of AI and showcase what a new imagination and its implementation in AI techniques look like to better recognize them, we demonstrate our process to construct a new imagination of “AI for just work” and apply it in our specific AI development. We first construct and justify three foundational assumptions that form the value system for the new imagination, and deduce new properties of AI from these assumptions. We then show the process of how the new imagination guides specific technical decisions in the case study of medical image synthesis task.
@@ 4.1 Foundational assumptions
We define complex phenomenon by borrowing the definition of complex systems, which are “co-evolving multilayer networks,” are context-dependent, and are composed of many non-linearly interacting elements [102], such as language, human behavior, human mind, a living cell, financial market, social or natural phenomena. A1 is justified by our counterarguments to S1 and the normal worldview of science and technology (Fig. 2-A). This worldview acknowledges the limits of science and technology that science “isn’t a window onto a disembodied, God’s-eye perspective. It doesn’t grant us access to a perfectly knowable, timeless objective reality. …Instead, all science is always our science, profoundly and irreducibly human, an expression of how we experience and interact with the world” [47]. Similarly, A1 assumes that there does not exist a disembodied, God-like intelligence of knowing that grants us access to a perfect world model that can replace the original phenomena. It acknowledges and respects the fact that AI models are built upon human experience (embodied as human labor, prior knowledge, dataset, infrastructure, etc.), natural resources and materials of our planet [38]. This assumption is also well-supported by common framing in the academic community interested in machine learning and statistics. However, it stands in contrast to how AI is sometimes described in the tech industry.
@@
A2 is justified by our counterarguments to S2 that the root causes of AI-based algorithmic oppression and exploitation are in the social hierarchy that wrongly assumes some people and their work are inferior, and in the imbalanced power between the oppressor and the oppressed. To correct the unjust assumptions, A2 assumes that workers and their work are not devalued or degraded by any means, including but not limited to: workers’ age, gender, sexuality, race, ethnicity, religion, nationality, physical or mental conditions, educational level, socioeconomic status, whether the work is paid or unpaid, the type, role, and salary of the work, the amount of efforts or skills required or engaged, how workers are rated and their ratings, etc. It means that all workers have equal social recognition, social respect, social support, visibility, power, autonomy in work, and reasonable salary that reflects a comparable level of respect (if the work is paid), and no workers and their work are worth more than any other workers and their work. Furthermore, in this new imagination, we refer to critical theory and adopt a broader definition of work to revive its original meaning as “meaningful, fulfilling activities” as the “social means necessary to live a flourishing life” [110]. Work under this assumption includes a wider range of human activities that may not count as “work” in the current narrowly framed economic system based on market exchange, such as reproductive labor and emotional labor [91, 77]. In the context of AI, workers include a wide range of people who work for or behind AI (such as data labor), who work with AI (such as users), and whose work or life is impacted by AI (such as stakeholders). Using the broad concept of work ensures that people and their visible or invisible labor and contributions can be recognized, supported, and deserve workers’ power to protect their rights. In the rest of the paper, we use the word “workers” to denote the broad groups of people that include workers, users, stakeholders, and the public. The ideas in A2 are also justified by AI ethics principles of autonomy and justice [45].
@@
A3 is justified by our counterarguments to S3 that merely pursuing productivity and efficiency cannot automatically lead to shared prosperity, as this goal can easily be hijacked by the dominant power to serve the benefits and interests of the powerful. Instead of pursuing progress and productivity at the expenses of essential ethical values, it is more reasonable to maintain dynamic balance between ethics and development for the same goal of the mutual flourishing of humanity and the planet. The ideas in A3 are also justified by AI ethics principles of beneficence and non-maleficence [45].
@@ 4.2 AI properties
As shown in A1 and Fig. 2, the blind spot worldview [47] and the abstraction nature of AI models tend to get people caught up in the upper abstraction space, ignoring the grounded real-world space where AI models are applied and can have an impact. As Rolnick et al., [95] point out, “ML algorithms designed in blue-sky, methods-focused research continue to fall short when used directly for applications.” According to A1, because AI models abstract and simplify real-world phenomena, the modeling process creates inevitable information loss due to the gap between concrete world and abstract model. Therefore, considerable human efforts are indispensable (visualized as pillars in Fig. 2-A) to mitigate the gap and information loss. Human effort acts in a two-way process from real world to AI model and back to real world: First, in the modeling process, human labor is needed to determine which information from the real world is relevant for AI abstraction, such as the efforts in problem formulation, data curation, data labeling, and model design; Second, to apply AI in the real world, human labor is necessary to act as a buffer between AI model and real-world tasks to deal with the limitations of model and the ambiguity and fluidity of the real world, such as efforts in model evaluation, reorganizing the existing workflow, incorporating AI outputs in real-world tasks, long-term model monitoring and quality control. The current main paradigm of methods-focused AI research is not driven by and grounded in the real world problems, making the above indispensable human/social processes not being seriously considered and incorporated in the modeling process. This greatly limits the positive impacts of AI research to society and impedes AI research itself [95]. Furthermore, failing to ground AI models in real world applications tends to undervalue human labor in AI development and deployment, and underestimate the negative impacts of AI implementation on workers, which violates A2. P1 joins the recurring call in the AI community for real world problem-grounded AI [95, 97, 104], and provides additional justifications for this call based on the new imagination.
@@
To ground the problem formulation, design, and evaluation of AI in real-world tasks, the first step is to understand and acknowledge the inevitable gap between AI modeling in the abstracted space and its application in the real world, and should not conflate the model’s capability at the abstracted level (confined by many idealized assumptions) with the model’s real-world utility and performance, which manifests the pathology of “confusing the map with the territory” in the blind spot worldview [47]. The second step is trying to mitigate the gap by relaxing the too idealistic assumptions confined by the toy problem formulation and benchmarks, and incorporating more domain- and task-specific knowledge and human experience in the modeling process. As Andrews et al., [13] state, this may improve scientific validity and avoid unethical pitfalls for AI. For example, instead of framing an evaluation problem as AI vs. doctor on a benchmark dataset, which is unrealistic in clinical scenario, the ultimate evaluation of a medical AI should assess how the AI and clinical workers perform when AI is embedded in clinical contexts [32].
@@
The current evaluation paradigm of AI is lopsided as it mainly evaluates the positive aspects of the proposed AI technique, such as performance improvement and efficacy of a proposed function from ablation study, but the negative aspects of the proposed AI technique, such as its limitations and negative impacts, are at most briefly mentioned without an equivalent amount of thorough analysis as the positive aspects [59]. The biased evaluation paradigm is driven by and reinforces the MIA that technical progress of AI is to reach an idealized image of AI that is perfect with no weaknesses; and technical progress is prioritized and harms can be fixed later (S1), which violates A3 in the new imagination. In the new imagination, we acknowledge that there does not exist a perfect state of AI, and AI always has its limitations and weaknesses no matter how technology progresses (A1). To avoid weakening workers’ power from the incomplete evaluation that only emphasizes AI strengths and human weaknesses but neglects human strengths and AI weaknesses (A2), and to avoid developing AI for the sake of technical progress and ignoring its negative aspects (A3), we propose that a thorough limitation analysis of AI should be a routine in AI evaluation. The limitation analysis includes the following aspects: 1) quantitatively and/or qualitatively assessing the scope, weaknesses, and failure modes, and associating potential social consequences with the limitations, so that users can have a holistic understanding of the strengths and risks of the proposed AI technique. 2) Being transparent and acknowledging the design-specific assumptions and limitations in problem formulation, model design, and evaluation methodology that may be mitigated in future work. 3) Acknowledging the intrinsic limits of the proposed AI technique that cannot be overcome with technological progresses. Setting up the thorough limitation analysis as an AI evaluation standard also aligns with the scientific practice in biomedicine or engineering that equally evaluate the scope, efficacy, side effects, and safety issues of a medication or an engineering product.
@@
Because AI models and applications have scopes, limitations, risks, costs, and uncertainty (A1, P2), it is reasonable to deduce that AI may not be applicable in every scenario, thus AI and its data collections can be rejected. As Benjamin, [16] states, informed refusal is the corollary of informed consent in ethics. Refusal also corresponds to the right to withdraw in research ethics, and has already been implemented in laws. For example, GDPR, [51] states “data subject shall have the right not to be subject to a decision based solely on automated processing,” and have the right to erasure data (“right to be forgotten”). These measures are meant to protect rights and shift power dynamics in favor of the powerless, in this innately imbalanced power structure where technical sectors and institutions are the dominant.
However, this refusal assumption is currently not embedded in AI technical development, which creates a strong framing effect that sets the given AI model and technical solutions as default and their necessity is rarely interrogated or challenged. As we mentioned previously in S4, not all problems are suitable to be solved by AI. Maybe the “problem” is not a real problem at all that needs to be solved, or maybe the root causes of the problem can be tackled by non-technical or low-technical approaches. Failing to recognize this framing effect falls into the “solutionism trap” [99]. Encoding refusal as the basic property of AI can avoid the trap of regarding AI and its data collection as the default and inevitable option. Furthermore, refusal is not the end goal, but a way of constructing a reciprocal relationship between the powerless and the powerful, such that the design of AI and data collection can incorporate goals and perspectives of workers, especially the vulnerable and marginalized, at the very start [16, 113], which is in line with A2, A3, and P1.
@@
Because certain AI design choices, such as automation, can create a displacement effect that reduces workers’ wages and weakens their power [6, 4], which devalues workers’ work thus violates A2, an ethical and prudent design of AI should take the impact of AI on worker rights and power into consideration, and choose to design AI that could improve, rather than weaken workers’ value in work. As we mentioned in the counterarguments to S3, such design choices can include AI that aids workers, AI that creates new labor-intensive tasks and opportunities in which workers have comparative advantages over machine [1]. In cases when AI displacement effect cannot be completely offset, mechanisms should be placed to compensate for the reduced workers’ rights and power, such as retraining workers for new skills. Scholars also point out that the so-called “collaborative AI,” “human-augmenting AI,” or “human-centered AI” may not necessarily lead to increased worker power, as some may still reproduce a hierarchical and exploitative relationship that shifts the oppression to the less powerful [14, 37]. Therefore, whether or not an AI design reduces worker power should not be determined by the powerful party of the technical sector or the institution who implements the AI, but should undergo democratic processes by incorporating a wide range of workers’ voices and perspectives in AI design, especially the vulnerable and marginalized. Ultimately, this collective design paradigm in AI development depends on power check mechanisms and structural changes that shift power dynamics from capital to workers and tackle the root causes in social institutions that exploit workers and make work miserable [49], such as collective bargaining and legislation to protect worker rights and power (for instance, their rights to reject AI in P3), and changing the ownership of production from private-owned to public- or worker-owned. The AI community has the agency to propose technical works and non-technical initiatives to facilitate, rather than impede, such power checks and structural changes.